# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@ Appendix E Additional Experimental Results
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? [Uncaptioned image] WRS {$S_{a}$: 3.96, $S_{g}$: 5.05, $S_{g}^{ni}$: 5.21, $S_{r}$: 5.34, $S_{p}^{ni}$: 5.38, $S_{c}$: 5.46, $S_{m}$: 5.75, $S_{p}$: 6.28, $S_{b}$: 6.9} {P2: 5.18, P1: 5.2, P3: 5.35, P0: 5.46} S with low statistical bias: $S_{a}$. S with high statistical bias: $S_{p}$. P that led to more statistical bias: P0 Analysis with more discrepancy: Inter-industry
@@
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? [Uncaptioned image] PIE % {$S_{g}^{ni}$: 523.09, $S_{g}$: 621.68, $S_{c}$: 742.19, $S_{a}$: 1206.44, $S_{m}$: 1466.13, $S_{p}^{ni}$: 1655.94, $S_{p}$: 2677.62, $S_{r}$: 2733.7, $S_{b}$: 5197.51} {P1: 995.19, P2: 1252.2, P3: 1563.94} S with low confounding bias: $S_{g}^{ni}$. S with high confounding bias: $S_{p}$. P that led to more confounding bias: P3. Confounder that led to more bias: Industry
@@
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? [Uncaptioned image] APE {$S_{g}^{ni}$: 6.49, $S_{g}$: 8.34, $S_{c}$: 8.42, $S_{m}$: 17.46, $S_{p}^{ni}$: 21.94, $S_{a}$: 28.28, $S_{p}$: 28.53, $S_{r}$: 29.65, $S_{b}$: 50.37} {P2: 12.74, P3: 17.34, P1: 21.11} S with low APE: $S_{g}^{ni}$. S with high APE: $S_{p}$. P with low APE: P2. P with high APE: P1. Confounder that led to high APE: Company
@@
RQ4: Does $P$ affect the accuracy of $S$? This hypothesis does not necessitate a causal model for its evaluation. SMAPE, MASE, Sign Accuracy SMAPE: {$S_{c}$: 0.05, $S_{a}$: 0.06, $S_{g}$: 0.06, $S_{g}^{ni}$: 0.07, $S_{p}^{ni}$: 0.09, $S_{p}$: 0.1, $S_{m}$: 0.1, $S_{r}$: 0.83, $S_{b}$: 1.28}; MASE: {$S_{c}$: 4.67, $S_{g}$: 5.04, $S_{g}^{ni}$: 6.75, $S_{p}^{ni}$: 7.91, $S_{a}$: 8.54, $S_{p}$: 9.03, $S_{m}$: 9.13, $S_{r}$: 86.74, $S_{b}$: 947.56}; Sign Accuracy: {$S_{m}$: 40.91, $S_{p}$: 44.42, $S_{p}^{ni}$: 47.4, $S_{r}$: 49.71, $S_{g}$: 50.93, $S_{c}$: 51.99, $S_{g}^{ni}$: 52.94, $S_{a}$: 58.57, $S_{b}$: 62.6} SMAPE: {P3: 0.06, P0: 0.07, P1: 0.08, P2: 0.08}; MASE: {P3: 6.11, P0: 6.87, P2: 7.22, P1: 8.99}; Sign Accuracy: {P2: 49.26, P1: 49.42, P3: 49.73, P0: 49.97} S with good performance: $S_{c}$. S with poor performance: $S_{m}$. P with high impact on performance: P2.
@@
Table 4: Summary of the research questions answered in the paper, causal diagram, metrics used in the experiment, average of the metric values compared across different systems, average computed across different perturbations, and the key conclusions drawn from the experiment. Overall, multi-modal FMTS demonstrated greater robustness and forecasting accuracy compared to multi-modal FMTS. TS FMTS demonstrated greater robustness and forecasting accuracy compared to GP FMTS. All the raw scores and ratings are shown in Table 2.
(a) SMAPE
(b) MASE
(c) Sign Accuracy
Figure 15: Bar plots showing the robustness metrics values across different systems and perturbations.
(a) WRS
(b) PIE % scores
(c) APE scores
Figure 16: Bar plots showing the robustness metrics values across different systems and perturbations.
Figure 17: Radar plots showing robustness metrics for all FMTS and $S_{a}$ under different perturbations.
Figure 18: Radar plots showing forecasting accuracy metrics for all systems under different perturbations.
Figure 19: Heatmap for all metrics for all models. Lighter shade indicates better performance.
Appendix F Additional User Study Results
In this section, we present all the hypotheses, the results from the statistical tests conducted to validate these hypotheses, and the conclusions drawn from the results.
@@ Appendix F Additional User Study Results
Table 5: Summary of one sample right-tailed t-test results: Comparison of sample means to the hypothesized mean of 2 with a sample size of 26. The right-tailed p-values indicate whether the sample means are significantly greater the hypothesized mean. denotes that mean of responses for all the questions is greater than 2.
@@
Table 6: Table with the hypotheses evaluated in the user study, statistical tests used to validate the hypotheses, results obtained, and conclusions drawn.
Appendix G Source Code for Data Processing
1# Convert data from Yahoo! finance to sliding window format.
2def sliding_window(data, window_size, company):
3 sequences = []
4 for i in range(len(data) - window_size):
5 seq = data[i:(i + window_size + 1)].tolist()
6 sequences.append([company] + seq)
7 return pd.DataFrame(sequences)
8
9# Perturbations:
10# Drop-to-zero: Every 80th stock price in the numerical data will be turned into zero.
11def drop_to_zero(df, col):
12
13 new_df = df.copy()
14 new_df.loc[new_df.index % 80 == 0, col] = 0
15
16 return new_df
17
18# Value halved: Every 80th stock price in the numerical data will be halved.
19def value_halved(df, col):
20
21 new_df = df.copy()
22 new_df.loc[new_df.index % 80 == 0, col] /= 2
23
24 return new_df
25
26# Missing values: Every 80th stock price in the numerical data will be NaN’.
27def missing_values(df, col):
28
29 new_df = df.copy()
30 new_df.loc[new_df.index % 80 == 0, col] = float(’nan’)
31
32 return new_df
33
34# Code to generate time series line plots.
35def plot_ts(input_path, output_path):
36 data = pd.read_csv(input_path)
37
38 companies = data.iloc[:, 0]
39 time_steps = data.iloc[:, 1:]
40
41 for i, company in enumerate(companies):
42 plt.figure(figsize=(12, 6))
43 plt.plot(time_steps.columns, time_steps.iloc[i], marker=’o’)
44 plt.title(fTime Series for {company}’, fontsize=19)
45 plt.xlabel(’Time Steps’, fontsize=17)
46 plt.ylabel(’Values’, fontsize=17)
47 plt.grid(True)
48 x_ticks = time_steps.columns[::5]
49 plt.xticks(x_ticks, rotation=45, fontsize=15)
50 plt.yticks(fontsize=15)
51 plt.tight_layout()

### v3

@@ 5.3. Overall Performance Comparison
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? [Uncaptioned image] WRS $S_{a}$: 3.96, $S_{g}$: 5.05, $S_{r}$: 5.15, $S_{v2}$: 5.20, $S_{g}^{ni}$: 5.44, $S_{c}$: 5.46, $S_{p}^{ni}$: 5.48, $S_{v1}$: 5.71, $S_{m}$: 5.75, $S_{p}$: 6.27, $S_{b}$: 6.9} {P4: 5.42, P1: 5.49, P3: 5.51, P5: 5.52, P6: 5.52, P2: 5.55, P0: 5.70} S with low statistical bias: $S_{a}$. S with high statistical bias: $S_{p}$. P that led to more statistical bias: P0 Analysis with more discrepancy: Inter-industry
@@
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? [Uncaptioned image] PIE % $S_{g}^{ni}$: 1107.66, $S_{g}$: 1115.08, $S_{v2}$: 1346.46, $S_{a}$: 1448.29, $S_{v1}$: 1848.20, $S_{p}$: 2459.30, $S_{m}$: 2544.20, $S_{r}$: 2668.52, $S_{c}$: 2755.50, $S_{p}^{ni}$: 2778.06, $S_{b}$: 4758.16} {P1: 1711.88, P4: 2035.95, P3: 2057.31, P6: 2410.28, P2: 2628.52, P5: 3646.20} S with low confounding bias: $S_{g}^{ni}$. S with high confounding bias: $S_{p}^{ni}$. P that led to more confounding bias: P5. Confounder that led to more bias: Industry
@@
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? [Uncaptioned image] APE $S_{g}^{ni}$: 5.89, $S_{v1}$: 7.34, $S_{c}$: 7.80, $S_{m}$: 9.83, $S_{g}$: 6.46, $S_{v1}$: 6.46, $S_{p}^{ni}$: 12.66, $S_{p}$: 15.98, $S_{r}$: 21.57, $S_{a}$: 27.73, $S_{b}$: 33.95} {P3: 9.96, P6: 12.2, P2: 12.9, P4: 13.19, P5: 18.36, P1: 20.25} S with low APE: $S_{g}^{ni}$. S with high APE: $S_{a}$. P with low APE: P3. P with high APE: P1. Confounder that led to high APE: Industry
@@
RQ4: Does $P$ affect the accuracy of $S$? This hypothesis does not necessitate a causal model for its evaluation. SMAPE, MASE, Sign Accuracy SMAPE: $S_{c}$: 0.051, $S_{v1}$: 0.053, $S_{g}$: 0.055, $S_{a}$: 0.058, $S_{v2}$: 0.06, $S_{g}^{ni}$: 0.06, $S_{r}$: 0.83, $S_{p}^{ni}$: 0.084, $S_{p}$: 0.097, $S_{m}$: 0.098, $S_{b}$: 1.276}; MASE: $S_{c}$: 4.67, $S_{v1}$: 4.80, $S_{g}$: 5.04, $S_{v2}$: 5.49, $S_{g}^{ni}$: 5.76, $S_{a}$: 8.54, $S_{p}^{ni}$: 7.60, $S_{p}$: 9.02, $S_{m}$: 9.13, $S_{r}$: 86.76, $S_{b}$: 947.56 }; Sign Accuracy: $S_{m}$: 40.91, $S_{p}$: 44.42, $S_{p}^{ni}$: 45.24, $S_{v2}$: 49.33, $S_{r}$: 49.75, $S_{g}$: 50.93, $S_{v1}$: 51.34, $S_{g}^{ni}$: 51.37, $S_{c}$: 51.99, $S_{a}$: 58.57, $S_{b}$: 62.6} SMAPE: {P0: 0.24, P2: 0.25, P1: 0.26, P3: 0.28, P4: 0.38, P5: 0.38, P6: 0.39}; MASE: {P0: 99.06, P2: 99.57, P1: 101.27, P3: 119.66, P4: 175.56, P5: 175.56, P6: 176.43}; Sign Accuracy: {P1: 49.86, P0: 51.35, P2: 51, P3: 51.16, P4: 51.79, P5: 51.43, P6: 50}; S with good performance: $S_{c}$. S with poor performance: $S_{m}$. P with high impact on performance: P6.
@@
Table 2. Summary of the research questions answered in the paper, causal diagram, metrics used in the experiment, average of the metric values compared across different systems, average computed across different perturbations, and the key conclusions drawn from the experiment. All the raw scores and ratings are shown in Tables 5 and 6.
Now, we provide an overall comparison of the different systems across all metrics to highlight key findings about their performance under various perturbations by answering the additional research questions (ARQs) stated in Section 1. Figures 5 and 6 show radar plots with robustness metrics and forecasting accuracy, respectively.
Clear Domination Signals: From Figure 5 and the detailed results from Section 5.2, we can draw the following conclusions:
$S_{c}$’s Superiority in Forecasting Metrics: $S_{c}$ consistently outperformed other models in terms of forecasting accuracy metrics, specifically SMAPE and MASE. This indicates that $S_{c}$ is highly effective in predicting stock prices with minimal error.
General Superiority Over Biased and Random Systems: All models perform better than the biased and random systems in forecasting metrics. This underscores the importance of using well-designed models over naive or biased approaches.
Robustness in PIE % and APE Metrics: According to the average scores, the $S_{g}^{ni}$ system demonstrated superior robustness in PIE % and APE metrics. This suggests that $S_{g}^{ni}$ is more resilient to perturbations and confounding biases compared to other systems.
Role of Confounders: Our analysis (Figs. 4 and 5) shows that using industry as a confounder introduces more bias, with higher PIE% and APE scores indicating significant industry-specific effects on the relationship between perturbations and residuals. Inter-industry comparisons also show more discrepancies, as evidenced by WRS scores.
Figure 7. Radar plots comparing the average robustness and accuracy metrics for (left) uni-modal vs. multi-modal models, and (right) foundation vs. non-foundation models. Lower values (i.e., lines closer to the center) indicate lower performance, while points farther from the center represent better performance across metrics.
ARQ1: How do different model types compare in performance? (Foundation vs. Non-foundation)
Answer: As shown in Figure 7, Non-foundation models (ViT-num-spec models) demonstrate superior performance over foundation models across both robustness and forecasting accuracy metrics.
ARQ2: Does multi-modality improve the performance of TSFM?
Answer: The results in Figure 8 show the effect of multi-modality on model performance. While $S_{g}$ (uni-modal) outperforms $S_{g}^{ni}$ (multi-modal) and $S_{p}^{ni}$ (multi-modal) outperforms $S_{p}$ (uni-modal) in terms of individual metrics, the overall average trend across all metrics suggests that incorporating multiple modalities can lead to more balanced and improved performance in both robustness and accuracy.
Figure 8. Effect of the modalities for $S_{g}$ (left) and $S_{p}$ (right).
ARQ3: How does the architecture of FMs influence performance? (Time-series-specific vs. general purpose and encoder-only vs. decoder-only vs. encoder-decoder)
Answer: Our evaluation (Fig. 9, left) indicates that the Time Series architecture generally performs better across several metrics, such as achieving the best values in APE_C, PIE_C, SMAPE, MASE, and WRS_I, suggesting that the TS architecture may be more effective for these specific tasks compared to the general purpose architectures. Fig. 9 (right) shows that decoder-only architecture outperforms others in terms of both accuracy and robustness.
Figure 9. Role of architecture in forecasting accuracy and robustness. Performance is averaged across models within each category. See Table 1.
Overall, our rating method highlights $S_{c}$’s forecasting accuracy, the robustness of multimodal systems against perturbations and confounding biases, and the superiority of well-designed models over naive approaches.
5.4. User Study
@@
Table 3. Summary of one sample right-tailed t-test results: Comparison of sample means to the hypothesized mean of 2 with a sample size of 26. The right-tailed p-values indicate whether the sample means are significantly greater the hypothesized mean. denotes that mean of responses for all the questions is greater than 2.
@@
Table 4. Table with the hypotheses evaluated in the user study, statistical tests used to validate the hypotheses, results obtained, and conclusions drawn.
We conducted a user study to evaluate the ratings generated by our approach for comparing the behavior of various TSFM based on two key metrics: robustness and statistical fairness (defined as lack of statistical bias). To simplify the evaluation for participants, we converted the generated ratings into rankings (i.e., the system with the highest robustness ranking is the most robust system). The main objective of this study was to validate the following hypotheses:
This IRB-approved study (IRB Exemption and Compensation: “This research study has been certifed as exempt from the IRB per 45 CFR 46.104(d)(3) and 45 CFR 46.111(a)(7) by the University of South Carolina IRB# Pro00138672 on 24/07/2024.”) involved participants evaluating FMTS models forecasting stock prices for companies in various industries, participants were introduced to key concepts including robustness, fairness, and error metrics (maximum residual, mean, and standard deviation of errors) to ensure informed evaluations. All inputs were sought on Likert 5-value scale. The details and complete set of questions can be found here: https://tinyurl.com/45u7hapn.
The study was structured into four panels: a self-assessment on knowledge about time series and financial data, a fairness panel, where fairness was the evaluated metric, and two robustness panels, where robustness was assessed under two different perturbations (P1 and P2). To make the plots and rankings easier to interpret for participants, we selected two perturbations (P1 and P2) from the six used in our experiments and six representative systems from the original set of eleven.
In the fairness panel, participants were presented with graphs depicting the residual values of six different systems and an ideal system using stock price data from the Technology and Pharmaceuticals industries. They were provided with the mean and standard deviation of errors and asked to rank the systems from least to most fair. Participants then rated the difficulty of this ranking task (1 being the most difficult). Subsequently, they were shown the rankings generated by our approach and asked to rate the accuracy of these rankings (1 being the least accurate). Finally, participants were asked to rank the difficulty of comparing the behavior of different systems using our rankings (1 being the most difficult).
For the robustness panels, similar questions were posed, with users evaluating systems based on their robustness to different perturbations (P1 and P2). A total of 26 users from academia and industry participated over two weeks. We performed different types of statistical tests (results are shown in Tables 3 and 4) to draw conclusions. We now discuss the key findings from the tests.